OpenStudy (anonymous):

Let f(x)=-5-4 and g(x)=6x-7. Find f(x) +g(x). a. -11x+3 b. x+3 c. -11x-11 d. x-11

OpenStudy (anonymous):

f(x) = -5-4?

OpenStudy (anonymous):

if f(x) = -5x-4 , then just add the two equations. -5x-4+6x-7 = x-11

OpenStudy (aziaboo45):

c

OpenStudy (aziaboo45):

the answer is c